40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 134     Entry time: Wed Nov 28 17:41:34 2007
Author: rob 
Type: Update 
Category: PSL 
Subject: FSS again 
I investigated the FSS a bit more today. I looked at the signals coming out of the FSS frequency reference, and saw that both the LO and PC drive were distorted, non-symmetric waveforms. In addition, the LO path had a 3dB attenuator, meaning the mixer was starved. I placed mini-circuits SLP-30 filters in both paths, and now both are nice sine waves. I also took out the 3dB att. With this work, and the CG slider maxed out at 30, the FSS open loop gain (for real this time) goes up to ~250kHz. Still needs more investigation.
ELOG V3.1.3-